# Numerical simulation of type II primordial black hole formation

### v5

@@ 3.1 Time evolution of type I/II fluctuations
First, we show the time evolutions of the fluid energy density and the lapse function in Fig. 3 for $\mu=0.5$, $1.2$, and $1.8$. For the cases of type I fluctuations $\mu=0.5$ (left) and $\mu=1.2$ (middle), the former disperses, but the latter contracts and forms a peak at the center in the density plot in Fig. 3. The latter case (middle) indicates a typical gravitational collapse, for which the lapse is decreasing to a very small value as plotted in Fig. 3 as Ref. [40]. For a type II fluctuation $\mu=1.8$ case (right), the collapsing behavior happens similarly to the $\mu=1.2$ (middle) case in the plots of density and lapse.
@@
Figure 3: Snapshots of the fluid energy density contrast $\rho/\rho_{\text{bg}}$ and lapse function $\alpha$ of type I fluctuations (left and middle) and a type II fluctuation (right). The initial configurations ($t=0.01t_{H}$) are shown in black on a different scale.
To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint555 First, we evaluate the appropriately normalized value of the constraint violation $\mathcal{H}_{p}$ at each grid point labeled by $p$. Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid points.  [21](Fig. 4). One can find roughly quadratic convergence in the different spatial resolutions, and it is consistent with our numerical code with a second-order scheme implemented in evaluating the fluid flux.
@@
Figure 4: The average value of the Hamiltonian constraint violation with different resolutions of the number of grid points for three typical cases for no collapse (left), collapse from a type I fluctuation (middle), and collapse from a type II fluctuation (right), respectively. The horizontal axis is a time coordinate normalized by the time of the horizon entry. We use the simulation result to $t=50t_{H}$ to evaluate spacetime structures around horizon formation before the violation propagation reaches the bulk region from the outer boundary. For evaluating a longer time evolution of the PBH mass, the average is taken outside the BH apparent horizon (or the future-outer trapping horizon, described in Sec. 3.2) after the horizon formation at around $t=40t_{H}$ for $\mu=1.2$ and $t=24t_{H}$ for $\mu=1.8$.
@@ 3.2 Identification of BH apparent horizon
First, let us introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$. We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$. A null expansion is expressed as $\theta\coloneqq 2\mathcal{L}_{l}R/R$, where $\mathcal{L}_{l}$ denotes the Lie derivative along $l$. Trapping horizons are defined by the surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $12M/R=$ 1 with $M$ being the Misner–Sharp mass defined by $M(t,r)\coloneqq R(1-\nabla_{\mu}R\,\nabla^{\mu}R)/2$ [29, 41, 42, 8].
@@
In Fig. 5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted. For $\mu=0.5$, the intersection between $\quantity(2M/R)(r)$ and the horizontal line of $12M/R=$ 1 is always only one. On the other hand, for the cases $\mu=1.2$ and $1.8$, the line intersects the $12M/R=$ 1 more than three times at a sufficiently late time. The worldlines of the intersections are shown in Fig. 6. To identify the future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs. 7 and 8. Comparing the directions of the null vector fields and contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig. 6 are future- and past-trapped regions, respectively. More specifically, in the future(past)-trapped region, both arrows direct to directions along which the areal radius decreases (increases), and the curve described by red points indicates the boundary satisfying $\theta\theta^{\prime}=0$. Then, we refer to a marginally future trapped surface that is the outer boundary of the future-trapped region on the $t=$const hypersurface as a BH apparent horizon. Similarly, we refer to a marginally past-trapped surface that is the inner boundary of the past-trapped region on the $t=$ const hypersurface as a cosmological apparent horizon.
@@
It is not easy to understand the horizon configuration. Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections. Readers who would like to get a quick intuition about the horizon configuration may refer to the schematic conformal diagram in Figures 16 and 17.
@@
Figure 5: Snapshots of the compactness $2M(t,r)/R(t,r)$. For the $\mu=1.2$ case, the trapping horizon at the center is formed at around $t=40t_{H}$ (middle). For the $\mu=1.8$ case, the horizon is formed at around $t=24t_{H}$ (right).
@@
Figure 6: Configurations of the trapping horizons $R(t,r)=2M(t,r)$. It corresponds to the points $(t,r)$ of the compactness $12M/R=$ 1 in Fig. 5.
Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.2$ case (type I) in the middle panel of Fig. 6. $R$-contour with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top is plotted as black curves.
Figure 8: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted as orange and blue arrows for $\mu=1.8$ case (type II) in the right panel of Fig. 6. The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top. In addition to the original plot of $R$-levels with 1.0 increments, it is also plotted with 0.1 increments from $R=8.1$ to $8.9$.
3.3 Spacetime structures of type I/II PBH
Let us introduce the classification of trapping horizons following Refs. [42, 43]. For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon. Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$. The classification of trapping horizons can be summarized as Table 1.
@@ 3.3 Spacetime structures of type I/II PBH
Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [42]. A black hole, white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with a future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
@@
Figure 9: The conformal diagrams of type A and B PBHs in radiation domination inferred from the present numerical results. The left figure defines a type A PBH as a causal structure of the typical case of PBH. The right one includes bifurcating trapping horizons and defines a type B PBH. The red and blue curves depict the future and past trapping horizons, respectively. The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type B case.
@@
Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions such as those shown in Figs. 7 and 8 for the radiation case. From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig. 9 as possible conformal diagrams of type I and type II PBHs, respectively.
@@
The crucial difference between the right and left panels in Fig. 9 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify PBH formation processes into types I and II; that is, we identify the PBH as of type I if there is no bifurcating trapping horizon and of type II if there appears at least one bifurcating trapping horizon during the time evolution associated with the intersection point. This difference has already been pointed out by KHW in Ref. [1] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [1] did not clearly distinguish the definitions of type I/II PBHs and type I/II fluctuations. However, in our simulation with $w=1/3$, we find that the horizon configuration is of type I for the $\mu=1.6$ case, but the initial fluctuation is of type II. Therefore, there is a gap between the definitions of type I/II PBHs and type I/II fluctuations. This fact is our motivation for the new definition of the PBH.
@@
The critical difference between the dust and the radiation cases is that trapping horizons in the FLRW regions are timelike and null for the former and the latter, respectively. See Figs. 15, 16 and 18 and Ref. [43]. Then, we confirm that the outermost boundary surface of the future-trapped region is the future-outer trapping horizon, which describes a BH apparent horizon. In addition, the outermost boundary surface between the past-trapped and non-trapped regions describes a cosmological apparent horizon, which asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
@@
It is worth noting that the schematic diagram given by Fig. 9 implies that all light rays emitted from the center reach not the future null infinity but the future singularity in the type II PBH. Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig. 8 seem consistent with this observation.
Precisely, one could also consider the critical case between type I and type II PBHs when the two bifurcating trapping horizons degenerate, then we additionally classify this case as a type IId PBH, which has just one bifurcating trapping horizon.

### v6

@@ 3.1 Time evolution of type I/II fluctuations
First, we show the time evolutions of the fluid energy density and the lapse function in Fig. 3 for $\mu=0.5$, $1.2$, and $1.8$. For the cases of type I fluctuations with $\mu=0.5$ (left) and $\mu=1.2$ (middle), the former disperses, but the latter contracts and forms a peak at the center in the density plot in Fig. 3. The latter case (middle) indicates a typical gravitational collapse, for which the lapse is decreasing to a very small value as plotted in Fig. 3 as in Ref. [PhysRevD.105.103538]. For a type II fluctuation case with $\mu=1.8$ (right), the collapsing behavior occurs similarly to the $\mu=1.2$ (middle) case in the density and lapse plots.
@@
Figure 3: Snapshots of the fluid energy density contrast $\rho/\rho_{\text{bg}}$ and lapse function $\alpha$ for $\zeta(r)$ given by Eq. (2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively. The initial configurations ($t=0.01t_{H}$) are shown at the top of each panel in black on a different scale. We can see that the evolution does not lead to black hole formation for $\mu=0.5$ but does for $\mu=1.2$ and $1.8$.
To check the convergence of the simulation, we plot the spatial average value of the Hamiltonian constraint [^5] First, we evaluate the appropriately normalized value of the constraint violation $\mathcal{H}_{p}$ at each grid point labeled by $p$. Then the average is taken as $\sum_{p=1}^{N}\mathcal{H}_{p}/N$, where $N$ is the total number of the grid.  [Harada:2015yda] (Fig. 4). One can find roughly quadratic convergence in the different spatial resolutions, and it is consistent with our numerical code with a second-order scheme implemented in evaluating the fluid flux.
@@
Figure 4: The average value of the Hamiltonian constraint violation with different resolutions of the number of grid points for $\zeta(r)$ given by Eq. (2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively. The horizontal axis is a time coordinate normalized by the time of the horizon entry. We use the simulation result upto $t=50t_{H}$ to deduce spacetime structures near the horizon formation before the violation propagation reaches the bulk region from the outer boundary. For computing a longer time evolution of the PBH mass, the average is taken outside the BH apparent horizon (or the future-outer trapping horizon, described in Sec. 3.2) after the horizon formation at around $t=40t_{H}$ for $\mu=1.2$ and $t=24t_{H}$ for $\mu=1.8$.
@@ 3.2 Identification of BH apparent horizon
First, we introduce two independent future-directed radial null vector fields, $l$ and $l^{\prime}$. We also define the null expansion for $l$ and $l^{\prime}$ as $\theta$ and $\theta^{\prime}$. A null expansion is expressed as $\theta\coloneqq 2\mathcal{L}_{l}R/R$, where $\mathcal{L}_{l}$ denotes the Lie derivative along $l$. A trapping horizon is defined by a hypersurface foliated by marginal surfaces that satisfy $\theta\theta^{\prime}=0$ or equivalently the compactness $2M/R=1$ with $M$ being the Misner–Sharp mass defined by
$M(t,r)\coloneqq\frac{1}{2}R(1-\nabla_{\mu}R\,\nabla^{\mu}R).$ (3.1)
A trapping horizon is said to be future, past and bifurcating for $\theta+\theta^{\prime}<0$, $\theta+\theta^{\prime}>0$ and $\theta+\theta^{\prime}=0$, respectively [PhysRev.136.B571, Hayward:1994bu, Hayward:1993wb, galaxies10060112, Maeda:2009tk]. A trapping horizon can be identified with an apparent horizon in the present setting. Black hole and cosmological apparent horizon are future and past trapping horizon, respectively.
In Fig. 5, the snapshots of the compactness $\quantity(2M/R)(z)$ are plotted for $\mu=0.5,1.2$ and $1.8$. For all these three cases, we can see that at the earliest time $t=0.01t_{H}$, the compactness monotonically increases from 0 at $z=0$ and crosses $1$ at some value of $z$, corresponding to a past trapping horizon or a cosmological horizon. For $\mu=0.5$, the compaction monotonically increases even in the later times and crosses $1$ only once even in the late times. This implies that no black hole forms in this case. For $\mu=1.2$, after some evolution, the monotonicity breaks down. It obtains a maximum and a minimum both smaller than $1$. The maximum is first smaller than $1$ and later gets larger than $1$, while the minimum continues to be below $1$. As the maximum exceeds $1$, the number of intersection points changes as $1$, 22, and $3$. This implies black hole formation inside the cosmological horizon. For $\mu=1.8$, at first sight, the behavior is similar to that of $\mu=1.2$, but we can see that the maximum and minimum are larger than $1$ for $t=6t_{H}$, when the monotonicity breaks down. As time proceeds further, the minimum decreases and gets smaller than $1$ for $t=12t_{H}$, while the maximum continues to be above $1$. Around $t=18t_{H}$, another set of maximum and minimum points appear between the pre-existing two extrema, and the minimum gets smaller than 1 for $t=24t_{H}$. The number of intersection points becomes five. Then the outer maximum decreases to 1 and increases again after touching the horizontal line of $2M/R=1$. As will be understood from Fig. 6, the future trapping horizon appears after this moment of the turnaround of the maximum. After that, a similar behavior is observed for the inner maximum. Then the minimum between the two maxima increases to a value larger than 1, and finally, there are 3 intersections with the line of $2M/R=1$.
@@
The worldlines of the intersection points, which correspond to trapping horizons, are shown in Fig. 6. We can see a single trapping horizon for $\mu=0.5$, two distinct trapping horizons for $\mu=1.2$ and two intersecting trapping horizons for $\mu=1.8$. To identify future-trapped ($\theta<0$ and $\theta^{\prime}<0$) and past-trapped ($\theta>0$ and $\theta^{\prime}>0$) regions, we also show directions of two independent null vector fields $l$ and $l^{\prime}$ together with the contour curves of the areal radius in Figs. 7 and 8. Comparing the directions of the null vector fields and the contour curves of the areal radius, we can realize that the red and blue shaded regions in Fig. 6 are future- and past-trapped regions, respectively. More specifically, in the future (past)-trapped region, the areal radius decreases (increases) along both of the arrows, and the red curve indicates the boundary satisfying $\theta\theta^{\prime}=0$. Then, the boundary of the future (past)-trapped region is a future (past) trapping horizon. We can see in Fig. 6 that there is only a single past trapping horizon for $\mu=0.5$ and separate future trapping horizon and past trapping horizon for $\mu=1.2$. For $\mu=1.8$, two trapping horizons intersect each other twice. The intersection points correspond to bifurcating trapping horizons. The enlarged views of Fig. 8 around the intersection points and the region surrounded by the red curves are provided in Fig. 9.
@@
It is not easy to understand the horizon configuration, in particular, for the case of $\mu=1.8$. Therefore, we carefully investigate it and present a more concrete classification of trapping horizons in the following sections. Readers who would like to get a quick intuition about the horizon configuration can refer to the schematic conformal diagram in Figs. 17 and 18.
@@
Figure 5: Snapshots of the compactness $2M(t,r)/R(t,r)$ for $\zeta(r)$ given by Eq. (2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively. For all these three cases, we can see that at the earliest time $t=0.01t_{H}$, the compactness monotonically increases from 0 at $z=0$ and crosses $1$ at some value of $z$, corresponding to a past trapping horizon or a cosmological horizon. For $\mu=0.5$, the compaction monotonically increases even in the later times and crosses $1$ only once even in the late times. This implies that no black hole forms in this case. For $\mu=1.2$, after some evolution, the monotonicity breaks down. It obtains a maximum and a minimum both smaller than $1$. The maximum is first smaller than $1$, becomes equal to 11 at around $t=40t_{H}$, and later gets larger than $1$, while the minimum continues to be below $1$. As the maximum exceeds $1$, the number of intersection points changes as $1$, 22, and $3$. This implies the formation of a black hole trapping horizon inside the cosmological horizon at around $t=40t_{H}$. For $\mu=1.8$, at first sight, the behavior is similar to that of $\mu=1.2$, but we can see that the maximum and minimum are larger than $1$ for $t=6t_{H}$ when the monotonicity breaks down. As time proceeds further, the minimum decreases and gets smaller than $1$ for $t=12t_{H}$, while the maximum continues to be above $1$. Around $t=18t_{H}$, another set of maximum and minimum points appear between the pre-existing extrema, and the minimum gets smaller than 1 for $t=24t_{H}$. As the zoomed subplot shows, the number of intersection points becomes five for $t=24t_{H}$. Then, the outer maximum decreases to 1 and increases again after touching the horizontal line of $2M/R=1$. As will be understood from Fig. 6, the future trapping horizon appears after this moment of the turnaround of the maximum. After that, a similar behavior is observed for the inner maximum. Then, the minimum between the two maxima increases to a value larger than 1, and finally, there are 3 intersections with the line of $2M/R=1$.
@@
Figure 6: Configurations of the trapping horizons $R(t,r)=2M(t,r)$ for $\zeta(r)$ given by Eq. (2.5) for $\mu=0.5$, $1.2$ and $1.8$ in the left, middle and right panels, respectively. It corresponds to the points $(t,r)$ of the compactness $2M/R=1$ in Fig. 5. We can see that there is only a single past trapping horizon for $\mu=0.5$ and separate future trapping horizon and past trapping horizon for $\mu=1.2$. For $\mu=1.8$, two trapping horizons intersect each other twice. The intersection points correspond to bifurcating trapping horizons. The region surrounded by the trapping horizons is untrapped with the null expansions $\theta>0$ and $\theta^{\prime}<0$. The future and past boundaries of this region are future and past trapping horizons with the ends being bifurcating trapping horizons, respectively.
Figure 7: Directions of two independent null vector fields $l$ and $l^{\prime}$ are depicted with blue and orange arrows in the top and bottom panels, respectively, for $\zeta(r)$ given by Eq. (2.5) for $\mu=1.2$. The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top. The red curves indicate trapping horizons satisfying $\theta\theta^{\prime}=0$. Comparing the directions of the null vector fields and the contour curves of the areal radius, we can see that the red and blue shaded regions in Fig. 6 are future- and past-trapped regions, respectively. More specifically, in the future (past)-trapped region, the areal radius decreases (increases) along both of the arrows.
Figure 8: Same as Fig. 7 but for $\mu=1.8$. The $R$ contour is plotted as black curves with 1.0 increments starting from $R=1.0$ and increasing from left bottom to right top. In addition to the original plot of $R$ levels with increments of 1.0, it is also plotted with increments of 0.1 from $R=8.1$ to $8.9$. We can see that two trapping horizons intersect each other twice. The intersection points correspond to bifurcating trapping horizons. Looking at these panels very carefully, we can see that the intersection points on the left-hand and right-hand sides are a maximum and a saddle, respectively, in terms of the areal radius. The region surrounded by the trapping horizons is untrapped with the null expansions $\theta>0$ and $\theta^{\prime}<0$. The future and past boundaries of this region are future and past trapping horizons with the ends being bifurcating trapping horizons, respectively. The enlarged views around the intersection points by the two red curves and the region surrounded by the red curves are provided in Fig. 9.
Figure 9: Same as Fig. 8 but the enlarged views of the complicated region near the intersection points. Note that the red curves divide the whole region into five, i.e., the bottom left (I), top middle (II), top right (III), bottom right (IV), and the region in the center of the panel surrounded by the red curves (V). On each panel, $R$ is small on the bottom left, top middle, and bottom right, while large on the top right. So, we can conclude that the left-hand and right-hand intersection points of the red curves correspond to a maximum and a saddle, respectively, in terms of $R$. We can see that the blue arrows on the top panel are tangent to the $R$-contour on the red curve convex upward, implying that $\theta=0$ there, while the orange arrows on the bottom panel are tangent to the $R$-contour on the red curve convex downward, implying that $\theta^{\prime}=0$. Thus, the signs of the null expansions $(\theta,\theta^{\prime})$ in regions I, II, III, and IV are given by (,+)(-,+), (,)(-,-), (,+)(-,+) and $(+,+)$, respectively. We can see that the signs are (+,)(+,-) in region V, implying that this region is untrapped. The part of the red curve convex upward that is the boundary of region V has the signs (0,)(0,-), implying that it is a future trapping horizon. The same analysis implies that the part of the red curve convex downward that is also the boundary of region V is a past trapping horizon.
3.3 Spacetime structures of type A/B PBH
Let us introduce the classification of trapping horizons following Refs. [Hayward:1993wb, Maeda:2009tk, Harada_2018]. For this purpose, to avoid confusion, we introduce another notation for the two radial null vector fields and associated expansions as $l_{\pm}$ and $\theta_{\pm}$ for a specific trapping horizon. Then we always set $\theta_{+}=0$ when considering a trapping horizon satisfying $\theta_{+}\theta_{-}=0$. The classification of the trapping horizons can be summarized in Table 1.
@@ 3.3 Spacetime structures of type A/B PBH
Table 1: Classification of trapping horizons satisfying $\theta_{+}\theta_{-}=0$ [Hayward:1993wb]. A black hole, a white hole, and cosmological trapping horizons with $w=1/3$ $(w=0)$ are identified with future-outer, past-outer, and past-degenerate (past-inner) trapping horizons, respectively.
@@
Figure 10: The conformal diagrams inferred from the present numerical results in radiation domination. The left panel defines a type A PBH as a causal structure of the typical case of PBH. The right one includes bifurcating trapping horizons and defines a type B PBH. The red and blue curves depict the future and past trapping horizons, respectively. The bifurcating trapping horizons ($\theta_{+}=\theta_{-}=0$) are depicted as circles in the type B case.
Unlike the dust case discussed in Appendix A, we can only obtain numerical solutions for radiation case, such as those shown in Figs. 7 and 8 for $\mu=1.2$ and $1.8$, respectively. The trajectories of $\theta=0$ and $\theta^{\prime}=0$ are independently given by two smooth curves indicated by red lines in the top and bottom panels, respectively. In the top panels, on the smooth red line in the bottom right, the blue arrows are tangent to the solid black lines indicating the contour of the areal radius, which implies $\theta=0$ there. In the bottom panels, on the other hand, the orange arrows are tangent to the solid black lines on the other smooth red line that is convex downward in the top left, which implies $\theta^{\prime}=0$ there. A very careful look at Figs. 8 and 9 reveals that the intersection points on the left-hand and right-hand sides are a maximum and a saddle, respectively, in terms of the areal radius. From these figures, we can find that the region surrounded by the trapping horizons, which we call region V, is untrapped with the null expansions $\theta>0$ and $\theta^{\prime}<0$. The part of the red curve convex upward that is the boundary of region V is a future trapping horizon, while the part of the red curve convex downward that is the boundary of region V is a past trapping horizon.
@@
From these figures and carefully comparing the similarities and differences from the exact conformal diagrams of the dust solution discussed in the Appendix A, we can deduce possible conformal diagrams and proper identifications of the trapping horizons for the radiation case as shown in the left and right panels in Fig. 10 as possible conformal diagrams of type A and type B PBHs, respectively.
@@
The crucial difference between the right and left panels in Fig. 10 is the existence of the bifurcating trapping horizons $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(0;*)$, which is described by the intersection points of the two trapping horizon trajectories in Fig. 8. In this paper, we propose to use this difference to classify the PBH formation processes into types A and B; that is, we identify the PBH as type A if there is no bifurcating trapping horizon and as type B if at least one bifurcating trapping horizon appears during the time evolution associated with the intersection point. KHW has already pointed out this difference in Ref. [PhysRevD.83.124025] for the case of the dust fluid. Unlike the current proposal, the authors in Ref. [PhysRevD.83.124025] did not distinguish between the definitions of type I/II and type A/B. However, in our simulation with $w=1/3$, we find the case of type II-A, the horizon configuration is of type A from type II fluctuation for the $\mu=1.4$ case. This fact is our motivation for the new definition of the PBH.
@@
The critical difference between the dust and the radiation cases is that trapping horizons in the FLRW regions are timelike and null for the former and the latter, respectively. See, e.g., Ref. [Harada_2018]. Then, we confirm that the outermost boundary surface of the future-trapped region is the future-outer trapping horizon, which describes a BH apparent horizon. Therefore, the outermost boundary surface between the past-trapped and untrapped regions, which describes a cosmological apparent horizon, asymptotically approaches the past degenerate trapping horizon in the background FLRW universe.
@@
It is worth noting that the schematic diagram given by Fig. 10 implies that any light ray emitted from the center reaches not the future null infinity but the future singularity of the type B PBH. Although we cannot make any general statement for the moment, the null trajectories indicated by the orange arrows in Fig. 8 seem consistent with this observation.